OpenStudy (anonymous):

Find The Next Two Numbers In The Pattern. 2,4,8,16,... A. 18, 20 B. 32,64 C. 20,24 D. 6,10

OpenStudy (dobby1):

b

OpenStudy (anonymous):

I think its c

OpenStudy (dobby1):

because 2*2=4 4*2=8 8*2=16 16*2=32 32*2=64

OpenStudy (anonymous):

yea shes right

OpenStudy (dobby1):

i am a he